OpenStudy (anonymous):

what is the equation of the line y= -0.25x +1 in standard form? Use interger coefficients

OpenStudy (texaschic101):

standard form : Ax + By = C y = -0.25x + 1 (add 0.25 to both sides) 0.25x + y = 1 (multiply the equation by 100 to get rid of the decimal) 25x + 100y = 100 you might better get a second opinion on this....

OpenStudy (texaschic101):

The reason I say that is because I don't know if you can solve it by turning the decimal into a fraction and solve it that way....let me show you.. .25 is the same as 1/4 y = -1/4x + 1 1/4x + y = 1 (multiply by 4 to get rid of fraction) x + 4y = 4 I am not sure which one is correct, or if you can do it either way.

OpenStudy (texaschic101):

@Hero ...can you help with this

OpenStudy (texaschic101):

@hartnn ....are both of these answers correct ?

hero (hero):

Looks good to me @texaschic101

OpenStudy (texaschic101):

so you can use either one ?

OpenStudy (texaschic101):

I think either one can be used. Whichever way you want to do it....by either multiplying by 100 or changing the decimal to a fraction

OpenStudy (anonymous):

thanks

hero (hero):

The smallest integer would likely be the correct one, so I would go with x + 4y = 4

hero (hero):

Because even if you got 25x + 100y = 100 You could still divide both sides by 25 to get x + 4y = 4

OpenStudy (anonymous):

Here's some alternative forms. 1.)y=-0.25(x-4) 2.)0.25x+y-1=0 Alternative form assuming x and y are real is. 3.)y=-0.25x+(1.+0.i)